Name: Thomas Santa Maria

Content and Structure – The paper

1. Has a clearly stated topic or question that it will address – 10

2. Does address the topic or question – 5

3. Has a clearly articulated and logical structure:

a. First paragraph lays out the topic or question – 10 b. Later paragraphs have topic sentences and clearly identifiable main ideas – 10 c. Transitions between paragraphs are logical – 10

4. Accurately presents facts and viewpoints of sources – 10

5. Provides sufficient textual evidence to support the argument – 10

6. Is concise, clear, and not too repetitive – 10

7. Has a clearly identifiable conclusion – 10

8. Has a convincing argument that draws together the evidence presented – 15

Total: 100 (A)

Style and English Usage – The paper

1. Is clear, direct, and appropriate in tone – 20

2. Is free of misspellings – 10

3. Is free of grammatical mistakes – 15

4. Avoids incomplete sentences – 10

5. Avoids run-on or overly complicated sentences – 10

6. Uses correct punctuation – 10

7. Is free of jargon, cliches, and misused words – 13

8. Cites quotations and/or references correctly – 10

Total: 98 (A)

Comments: Tom, just a few minor comments about what is a very good essay for the most part. First, it sounds somewhat odd to say that Henrietta’s story is “triumphant” (as you do in the second sentence). The triumphant aspect is the story of the scientific advances that have resulted from the HeLa cells’ availability for research. But that is in some ways only tangentially part of Henrietta’s own story.

In the second-to-last line of the first page you say the book is a “novel.” That is not correct, of course. This is a non-fiction story.